### v12

# Intergroup violence in bursts

Abstract
During intergroup confrontations, agitating stimuli such as opponents’ threats and provocations can trigger collective violence, even without the usual mechanisms of ingroup cooperation, such as norms with sanctions. We examine video recordings of street fights between groups of young men. Their violence sometimes breaks out in a burst, wherein a majority of participants starts fighting almost simultaneously. At other times, only few group members participate and it takes them more time to do so. This difference in commencing collective violence can be understood by perceiving it as a collective action dilemma. We adapt an Ising model to show that the proportion of group members who cannot or do not want to contribute to the public good—victory over opponents—predicts whether violence takes the form of a burst or not.
1 Introduction
In studies of intergroup conflict, the large numbers of casualties of wars between and within countries draw most attention [18, 64]. Yet, the most frequent occurrences of intergroup violence involve small groups, including subgroups of larger groups. We focus on small groups, and want to explain how violence committed by unarmed, non-professionals starts off. When observing pertinent videos, one often notices bursty starts of violence, where the majority of a group engages almost simultaneously in violence against their opponents. Before punches are thrown, individuals in the focal group and their opponents insult and threaten, and thereby agitate, one another. For this to result in collective violence by focal group members, a critical level of agitation has to be reached. A necessary condition is that focal group members are in close proximity and maintain face-to-face contact [20, 17, 49].
In one of the most influential studies of violence, Randall Collins [20] argues that antagonists need to circumvent an emotional barrier of tension and fear in order to be able to use violence. Surmounting the barrier requires that assailants attain a state of emotional dominance. In face-to-face interpersonal conflicts, they can reach emotional dominance when the other party appears vulnerable or weaker, or when the focal group is supported by an audience [20]. Collins’ notion of “forward panics” corresponds to the bursts we observed. In forward panics, assailants seize the opportunity to forcefully attack opponents in moments of weakness or vulnerability, for example when being isolated or falling down to the ground. This emotional barrier is meshed with another barrier known as the dilemma of collective action [58, 59], which has gained less attention in scholarly work on violence. We will present a formal model that explicates the conditions for groups to burst into collective violent action. At the core of the model is opponents’ agitation.
When agitation passes a critical level (for example, when one individual is pushed by an opponent), violence can break out in a burst. Alternatively, the confrontation can escalate more slowly and with fewer participants. Understanding how collective violence starts is complicated by the presence of other people, who may form an audience or try to de-escalate [42, 62, 84]. De-escalation happens often in street fights [63]. Given the conditions of proximity, agitation, and de-escalation, our goal is to explain bursts versus non-bursts. To this end we use the centennial Ising model from physics [46, 75]. It resulted in a Nobel Prize for Parisi (in 2021; [60]) and has been used in various studies of social influence [15, 74, 37, 29]. We adapt the model in a novel way, use it as an agent based model, and apply it to videos of street fights between groups of (mostly) young men.
2 Theory
Collective violence involves a dilemma of collective action [59, 58], so before predicting bursts, we must first address how the Ising model explains overcoming such dilemmas. For an unarmed focal group, the ingroup ties that matter are face-to-face contacts in close proximity [20]. We already know from public goods experiments that if participants regard a public good to be valuable, more than half of the participants are conditional cooperators willing to contribute if many others do [16]. In other words, conditional cooperators conform to their (weighted) average neighbor in the network [7, 78]. In our case, the public good is victory over opponents, which means that opponents flee or are wrestled to the ground. If victory is not feasible, a second tier public good is the prevention of ingroup humiliation. When confronted with opponents, most individuals experience uncertainty, which increases the likelihood of their conforming to the group [86, 53]. Conformity makes sense from an evolutionary perspective when payoffs are hard to predict [80]. Therefore, individuals’ motivation not only depends on their appraisal of the public good, but also their desire to align with their group members.
Now think of two individuals, each with the behavioral options to defect or cooperate. Behavioral options correspond to magnetic spins in the original Ising model, but that is not relevant here; we apply the model differently and do not lean on an analogy with the original. The possibilities are: (1) both individuals defect, which avoids exploitation but does not yield any public good; (2) one individual cooperates, yielding half of the public good, but is exploited by the freeriding other; or (3) both cooperate at a cost, which maximizes the public good. The dashed line in Fig. 1 plots these possibilities, with the normalized number of cooperators ($N_{C}/n$) from left to right. At the beginning, both defect (the local minimum of $H$ on the left). The vertical axis ($H$, or energy in the original) conveys that moving uphill (i.e., that one of the two will start contributing to the public good) is unlikely. After all, most individuals understand that fighting alone against a group is very dangerous. By contrast, if moving to lower levels of $H$ is possible, it will certainly happen; thus, if one of the two cooperates, the other conditional cooperator will join in, shifting $H$ to its global minimum on the right. The hill is the graphical representation of the dilemma of collective action; it is also drawn for a group of five. It shows that without a motivating force, it is very unlikely that a group of non-fighting individuals ($n\geq 2$) will overcome the hill and fight collectively, and collective action is even less likely in larger groups.
Figure 1: The dilemma of collective action presented as a hill between full defection (left) and full cooperation (right), with the proportion of cooperators $(N_{C}/n)$ on the horizontal axis. Data points are based on Equation 1, for $C$ = 1 and $D$ = 1/2. The vertical axis $(H/n)$ could be intuited as a negative likelihood; moving uphill from a state of defection $(N_{C}/n=0)$ is very unlikely. One line (dashed) is drawn for a dyad and one for a clique of five individuals. The larger the group is, the more rounded the hill becomes.
Nonetheless, some individuals may start fighting, which in turn may prompt others to join. When agitating stimuli, such as opponents’ threats, increase, focal group members become aroused and angry. Consequently, one or few individuals may accidentally cooperate even when most others do not (yet). This phenomenon of accidental or spontaneous cooperation is called “trembling hands” in game theory [23]. Among defectors, a few trembling hands can influence proximate others and may initiate a cascade of cooperation, or burst if the cascade proceeds rapidly. Among cooperators, trembling hands can have the opposite effect.
For the model, we define turmoil ($T$, or temperature in the original) as opponents’ behavior that agitates members of the focal group. Agitation arouses them; they produce the hormones adrenaline and cortisol that linger on for a while. In the model, opponents’ turmoil has an effect on the focal group at the aggregate level and affects everyone equally. In larger groups than ours, individuals standing closer to opponents are exposed to more turmoil than others standing further away. Locally varying turmoil in the Ising model as well as network variations are examined elsewhere [10]; here, we make the model not more complicated than necessary for our empirical study. The model demonstrates that if turmoil increases, there is no gradual increase of accidental cooperators. Instead, when turmoil reaches a critical level, $T_{c}$, fighting breaks out in a burst; see Figure 2. Bursts clearly differ from a gradual build up of collective action and the distinction does not follow from other models in the literature. One might expect, by contrast, that the tipping point is smooth rather than sudden (it is a second order phase transition; [55]), and in large random networks (e.g., $n=1000$) it is. However, all large social networks are decomposed into smaller clusters, where the transition is sudden rather than smooth, also in time (Fig. 5), and in small clusters, the tipping point is reached first [10].
The distinction between bursty and slowly increasing collective action turns out to depend on the critical proportion of unconditional defectors (versus conditional cooperators) in the focal group, $p_{c}$. Perhaps they would be conditional cooperators under different circumstances, but they were not in our study. Hence, we use $p_{c}$ as a prediction that we test empirically. In violent situations, many people do not fight, for all kinds of reasons. They may be too scared to fight [20], have empathy with their opponents, disagree with violence (i.e., appraise the public good differently), feel no solidarity with their group [19], try to de-escalate, be stopped in their tracks by de-escalators (also called guardians in criminological studies), or they may have fought but were wounded, were wrestled to the ground, or became exhausted at some point and became passive. In the population at large, preferences for violence will have a distribution [26] with a negative mean value that expresses a general dislike of violence. Only few individuals in the tail of the distribution will self-select into violent groups and value the public good positively. [^1] Milliff [51] suggested that these people perceive their situation as uncertain but controllable. If the proportion of unconditional defectors in these groups stays below the critical level ($p\leq p_{c}$), turmoil ($T>T_{c}$) will be followed by a burst of violence. Above the critical level ($p>p_{c}$), however, there is neither a critical level of turmoil nor a burst (Fig. 2). Note that $p$ incorporates not only de-escalators in the focal group, but also focal group members who have been prevented from fighting by de-escalators in the focal group or by others.
Figure 2: Simulation of Equation 1 over a range of turmoil ($T$) on the horizontal axis. The proportion of cooperators ($N_{C}/n$) in a fully connected network in the size range of our empirical study ($n=5$) is on the vertical axis; $C=1$ and $D=1/2$. The black line (top) depicts a burst in the group without unconditional defectors ($p=0$). The dashed blue line (bottom) depicts a gradual increase of cooperation in the group with two unconditional defectors ($p>p_{c}$).
Most models of collective action in the literature revolve around individual rewards and punishments, called selective incentives [58], on top of a share of the public good. These incentives require norms about (in)appropriate behavior in a given situation [25], as well as monitoring of group members [66], and transmission of information (i.e., gossip) through the group’s network, which leads to reputations [56] that feedback through selective incentives, with or without leaders. This package of mechanisms is crucial for ongoing cooperation in the longer run but needs time to develop, which may not be available when threats are imminent. The more time people have, the better they can prepare themselves, which is especially important for high-risk situations. Examples of well-prepared groups are police, soldiers, firefighters, and combat medics who receive professional training that enables them to cooperate effectively and respond to situational stimuli in predetermined manners rather than spontaneously. By contrast, ordinary citizens and amateur fighters lack training and team coordination. For non-professional groups, such as ours, the uncertainties of outcomes, benefits, and costs are higher.
Several earlier models of cooperation do not feature the usual package of cooperation mechanisms; namely, models of thresholds [33], cascades [82], and critical mass [48]. Therefore, these models are potentially useful for our study. They draw on the assumption that some people take the initiative or leaders set cooperation in motion. Yet in many cases, violence occurs in the absence of leaders [35]. Hence, an explanation of violence should not depend on their presence. The Ising model points out how cooperation can start spontaneously without leaders, galvanized by accidental cooperators rather than exceptionally zealous ones. If there are initiative takers or leaders in favor of violence [31], however, they can be accommodated in our model.
These earlier models also draw on strong rationality assumptions, such as perfect information about the numbers of cooperators at every moment. Yet, in violent and other uncertain situations, the degree of rationality is bounded due to incomplete or absent information about opponents and group members beyond face-to-face contact. Moreover, individuals often respond to provocations in ways that harm their interests, by overestimating their own abilities and underestimating the tenacity of their opponents. The Ising model is more parsimonious than the aforementioned models because it does not rely on assumptions of strong rationality. For the model it is enough if individuals find the public good valuable, without assuming how (in)accurate their expectations are. The model combines the precision of game theory, including a mapping on payoffs—usually unknown to the participants—with the behavioral spontaneity of chaotic and uncertain situations [76]. Beyond threshold models, this uncertainty have been assessed qualitatively [73] instead of modeling it precisely [45]. Uncertainty has been added to threshold models, too, which makes it possible to loosen the rationality assumptions. Then, a small probability, $\epsilon$, that an individual starts cooperating by random chance suggests that the probability that someone in a group cooperates increases exponentially with group size, $1-(1-\epsilon)^{n}$ [47]. However, this argument omits the interdependence of group members. Consequently, it contradicts our empirical finding that cooperation starts earlier in smaller groups, as predicted by the Ising model. Finally, whereas game theoretic approaches to conflict readily become complicated, even if there are only two individuals [69], the simplicity of the Ising model makes it possible to investigate large networks computationally.
Our work contributes to Collins’ [20] influential qualitative model of violence. One of its important contributions was that it shifted researchers’ attention to what actually happens in fights, for which they started using visual data. At the core, his model has potentially quantifiable variables capturing emotions, such as confrontational tension (agitation in our model), forward panic (our burst), and emotional dominance, to predict the beginning of violence and the dominating party. Empirical studies using these variables have remained qualitative [8, 34, 39], to which we contribute quantitative insights. We thus build on Collins’ work, also use visual data, but our general strategy is to use behavioral variables that are relatively easily observable. We minimize the number of variables from the literature that are hard to observe, such as reputations, norms, and solidarity, without downplaying their importance; they can be incorporated in an extended version of the Ising model [9] but we lack video recorded data to include them in our study. For example, if a pro-social norm is important and has been measured, it can be implemented as an external field (see below). Then, it will precipitate cooperation. Complementing Collins [20], our approach adds precision though modeling and explains collective action.
3 Model
Members of a (possibly fledgling) focal group, indexed $i$ or $j$, can defect, $D$, or contribute, $C$, to a public good, with $0<D<C$. Behavioral variable $S_{i}$ of conditional cooperators can take the value $S_{i}=C$ or $S_{i}=-D$, whereas unconditional defectors stay put at $S_{j}=-D$. Before a collective action, everyone in a focal group defects. Network ties among focal group members, $A_{ij}>0$, mean that $i$ is in close proximity to, and senses the behavior of, group member $j$, or else $A_{ij}=0$. Because people tend to respond to proportions of their social environment rather than absolute numbers [82, 28], ties are row-normalized, with $w_{ij}=A_{ij}/\sum_{j=1}^{n}A_{ij}$ such that $\sum_{j=1}^{n}w_{ij}=1$. Based on our video data, we assume that in our small groups ($n<10$), sensing is reciprocal at least to some degree, but not necessarily symmetrical, and that everyone senses all others, thus the network is fully connected. [^2] Temporary exceptions to full connectedness in our data were groups where some people’s view was blocked by objects, opponents, or bystanders, as well as one larger group ($n=14$). These cases we investigated through simulations. The Ising model is the following Hamiltonian equation [5, 81]
$H=-\sum_{i\neq j}^{n}w_{ij}S_{i}S_{j}.$ (1)
We do not assume that individuals know their payoffs in advance, but they will heuristically—and perhaps wrongly—distinguish between valuable ($C>D$) and non-valuable ($C<D$) public goods or, to the same effect, between potentially efficacious and non-efficacious actions [67]. Note that payoffs are not used in the model’s calculations but are defined to provide a meaningful interpretation. When $i$ chooses between $C$ and $D$ amid $N_{C}$ cooperators, payoffs for cooperation and defection are, respectively
$P_{C}$ $=\theta(N_{C}+1)/n-1$ (2)
$P_{D}$ $=\theta N_{C}/n+Q$ (3)
[…]
For our empirical study, we have to choose values for $C$ and $D$ to predict the critical level of unconditional defectors, $p_{c}$. The most obvious choice is $C=1$, as in game theory and the original Ising model. For $D$ we want to avoid two trivial values: if $D=0$, there is no dilemma (but a downward slope to the right in Fig. 1), and if $D=1$, there is no point in cooperating, which is equally valuable as defecting. Choosing $D$ to be maximally distant from the two trivial values seems to be a reasonable first approximation. Hence, we set $S=\{1,-1/2\}$ for all conditional cooperators, also in the examples [11]. For unconditional defectors $j$ we set $S_{j}=-1/2$, irrespective of their reasons. Note that all earlier Ising models had either the values $\{1,-1\}$ [15, 74, 29, 85] or $\{1,0\}$ [21]. Of all these models, only one represents a public goods game, in this case for two individuals [1, 68] whereas our model is applicable to groups of any size.
Beyond our empirical study, the payoffs in the asymmetric Ising model can be generalized by relating $C$ and $D$ to the symmetric model through a mapping $\{C,-D\}\rightarrow\{S_{0}+\Delta,S_{0}-\Delta\}$, with a bias $S_{0}=(C-D)/2$ with respect to 0, and the two behavioral options symmetrical at each side of $S_{0}$ at an offset $\Delta=\pm(C+D)/2$. [^4] It can be shown that the asymmetry in $S$ is equivalent to the symmetric model with an external field $2S_{0}$ [11]. In the payoffs, the bias and offset are expressed through $R=S_{0}/\Delta$. If $\Delta$ is set to a fixed value (here, 0.75), decreasing $S_{0}$ (here, 0.25) makes the public good and cooperation for it less valuable. This decline is equivalent to an increasing proportion of cooperating network neighbors winning over an actor to cooperate, also in other binary decision models [82, 33]. Increasing $S_{0}$ makes cooperation more valuable, and it corresponds to a decreasing proportion of cooperating network neighbors. If there are initiative takers, indexed $j$, they will have a higher $S_{0,j}$ than the other group members, which lowers $T_{c}$ in small networks.
Solving the Ising model boils down to minimizing $H$, which can be done analytically by a mean field approach (in the Supplementary information). We prove that, if $C=1$ and $D=1/2$, $p_{c}=1/3$. For inhomogeneous (such as clustered) networks, the mean field approach is inaccurate, but $H$ can also be minimized computationally, which is simpler and can deal with inhomogeneous networks.
The computational approach over a certain range of $T$ is as follows (Fig. 3). For a given level of turmoil, a network node $i$ is randomly picked and $H$ is calculated. For comparison, $i$’s current behavior is flipped, from $D$ to $C$ (or the other way around if $i$ cooperates), and $H^{\prime}$ with the flip is calculated. The flip is accepted and implemented if $H^{\prime}<H$ or with a certain chance that increases with $T$. In other words, increasing $T$ increases the amount of randomness (due to situational turmoil) in $i$’s decision. A behavioral change of $i$ affects others in $i$’s neighborhood when it is their turn to decide. Consecutive decisions are Monte Carlo steps in the Metropolis algorithm [5] that loops through great many Monte Carlo steps (counted by $t$ in Fig. 3) in order to allow individuals’ interdependent behavior to settle down. This procedure is repeated at increments of $T$, in our study with step size 0.01.
Figure 3: The Metropolis algorithm. Index $i$ runs over the conditional cooperators, not the unconditional defectors; $t$ is a counter of Monte Carlo steps; and $c_{r}$ is a random number in the range $0\leq c_{r}\leq 1$.
To illustrate the emergence of cooperation with increasing turmoil in a group of $n=5$, we apply the Metropolis algorithm; see Figure 2. First, we investigate the case without unconditional defectors. At low turmoil, collective action does not start (continuous line), but at a critical level $T_{c}$, (almost) everybody bursts into cooperation, with a maximum (where $N_{C}/n\approx 1$) close to $T_{c}$. Then, a small number of accidental cooperators wins over most others to join the collective action. The effect of turmoil is nonmonotonic and the level of cooperation decreases if $T$ keeps increasing beyond $T_{c}$, which means that extraordinarily strong turmoil becomes more confusing than motivating to fight.
If there are unconditional defectors, $T_{c}$ increases, which is more pronounced in larger networks, and maximum cooperation is lower than in the previous case because the number of conditional cooperators is lower. [^5] In Fig. 1, a proportion of unconditional defectors, $p$, means that no matter how many conditional cooperators contribute, $N_{C}/n\leq 1-p$. If the proportion of unconditional defectors surpasses a critical level, $p_{c}$, and $T$ increases above $T_{c}$, there is no burst (dashed blue line in Fig. 2) and cooperation increases gradually to a low level. In the mean field analysis, there is no (burst of) cooperation if $p_{c}>S_{0}/\Delta$, independent of network size and density. In simulations, $p_{c}$ is proportionally less precise in smaller networks (Table S3), but because our networks are complete, we stick to the mean field prediction. Based on the mean field analysis ($p_{c}=1/3$) and a nearly identical result from the simulations ($p_{c}=0.34$), [^6] If unconditional defectors are clustered together in a sparse network, they are less in the way of collective action of the remainder network (thus $p_{c}$ is higher) than if they are evenly spread out across the network. our main prediction is that $p_{c}=1/3$.
The critical threshold of turmoil increases with network size at a decreasing rate [10], but it also increases with the proportion of unconditional defectors. At $T_{c}$, simulations point out that cooperation starts in small clusters of conditional cooperators, which we will also test empirically. The onset in small groups is puzzling because larger groups have a better chance to win at lower individual costs. Yet if in a small (sub)group, someone starts fighting, he (and rarely she) accounts for a relatively sizeable proportion of his neighbors’ social contacts, and more readily wins them over to fight than in a large group where he would comprise a small minority. Fighting ends when exhaustion sets in, one party dominates (i.e., opponents flee or are wrestled to the ground), or others intervene.
4 Data and Methods
In studying violence, lab experiments lack the turmoil and emotional intensity of violent confrontations due to their obligation to meet ethical standards. Empirical field studies, in contrast, cannot be based on a random sample of participants or groups, yet they are invaluable to obtain a realistic view of violence [4]. We obtained 42 videos from websites such as YouTube, LiveLeak, and WorldStarHipHop using search terms with the English keywords “brawl,” “street fight,” and “assault.” This sample is not random with respect to violence, but it is random with respect to temporal unfolding and (sub)group size. Of these clips, 36 are from English-speaking countries (mainly the US and the UK, with one from Canada and one from India); five of the remaining clips are from the Netherlands, and one is from Colombia. We did not observe differences in relevant behavior related to the location of the recording. To keep distracting factors away from our analysis, we excluded clips with professional fighters, long range weapons, protective clothing, a referee, ambush attacks, or youths in a school yard. Most of our videos are recorded on phones by bystanders and are left-truncated; in all likelihood, there would have already been some turmoil that motivated bystanders to start filming. The shortest lasted 30 sec. and the longest was nearly 5 minutes (mean 101 sec.; s.d. 59 sec.). Out of a potential 2 x 42 groups, where the opponents in one analysis become the focal group in the next, 25 groups attacked a single individual rather than a group. Because a lone individual is unable to act collectively, this leaves 59 groups to examine; see the Supplementary information for all case studies and elaborate examples of interpreting and coding. One group had 14 members, but all other groups were small, $2\leq n<10$ (mean 3.6). They were simulated as cliques wherein everyone could sense one another unless there were obstacles or de-escalators obstructing contact.
The videos were coded using Noldus Observer XT 14 software. Clips were played at half speed many times over, and one of us discussed the coding of each with one or two assistants. The assistants were unaware of the theoretical expectations. Each of the 406 individuals was coded for group belonging, and their behavior was interpreted and coded on the timeline.
We coded violence when force was used against another’s body (punching, slapping, kicking, hitting, stomping) and/or when another person’s body was forcefully moved. People may defect (not use violence) for all kinds of reasons and due to various causes. To contrast everyone not fighting from conditional cooperators, we coded them as unconditional defectors if others were fighting while they were not. Among them we include de-escalators, as well as group members who were prevented to fight by de-escalators (who in turn might belong to either group or none).
We subsumed the following behaviors of members of the opponent group under turmoil for the focal group: aggressing (including fighting gestures); pulling off clothing (jackets or vests); pulling up pants (signaling readiness to fight); pointing toward opponents; provocative gesturing with fingers or hands (as an invitation to engage); bending forward toward an opponent; approaching the focal group; encroaching (invading opponents’ personal space through using or damaging objects belonging to them); teasing, such as lightly hitting or ridiculing; and violence. Because stumbling and falling signal vulnerability of opponents, which tends to agitate focal group members [20, 54, 83], we included these mishaps in our measure. It is likely that over relatively brief time intervals, the effects of turmoil accumulate. Therefore we calculated the level of turmoil from the beginning of the video until a focal group’s (first) maximum participation in violence. We did this by multiplying the duration of each instance of turmoil by the number of individuals involved, and adding up all these weighted instances.
Given the distraction caused by turmoil, it is not feasible for all group members to react within 1 second to a group member who initiates violence (as they might in a well-organized sports team), whereas 3 seconds is too long for the notion of burst to apply to our small groups. Hence we defined a burst as an outbreak of violence by at least half of the group (Fig. 2), or both individuals in a dyad, if they started fighting less than 2 seconds after the first, with a 5% margin. [^7] The requirement that $N_{C}\geq 0.5n$ is based on simulations of small networks just above $p_{c}$, at this proportion in large networks (e.g., $n=1000$), $N_{C}<0.5n$.
Our data do not enable us to assess causality. Instead, we assess if the patterns in the data support or refute the theoretical predictions.
4.1 Ethics
The use of videos for research purposes poses distinct ethical challenges, largely due to the non-anonymous content of the videos. However, ethical guidelines for digital spaces tend to be less restrictive [52], with the consent of the participants being less stringent for data acquired from the public domain, including the internet. While our video corpus is open for use and inspection by other researchers upon request, we require that they take the same measures to ensure the anonymity of the persons portrayed as we did.
5 Results
Of the 59 groups considered, there were 23 groups where violence started in a burst, 15 groups where violence was collective without a burst, and 21 cases of violence by a single group member. Turmoil preceded all collective violence with one exception, where two individuals suddenly assaulted a passive victim. The critical level of turmoil ($T_{c}$) for bursts is case-specific and depends on group size, both in absolute number and relative to the size of the opponent group, and on the proportion of unconditional defectors. Additionally, the use of weapons has an intimidating effect.
We confirm the finding of an earlier study [20] that fighting tends to start in small groups or in small subgroups of larger groups. As predicted [10], we found that small (sub)groups start fighting at lower levels of turmoil than larger groups, and lower levels are of course reached earlier. The local emergence (versus central coordination) of collective action in small groups has also been observed in protests [76]. The size-turmoil relationship is slightly disturbed by dyads more often facing a larger opponent group and being therefore less likely to fight collectively than triads. If there was one party that dominated, it was always the larger group, with only one exception, in line with earlier models of warfare when controlling for weapons [41].
In our data, bursts developed in 13 (37%) of the 35 smallest groups (dyads and triads) and in 11 (46%) of the 24 larger groups. In bursts, the correlation [^8] The predicted relation is $T_{c}\approx n^{1/2}$ [10], but due to the small data and unexplained variance therein, a curve does not fit better than a straight line; hence, the reported correlation. between focal group size and opponents’ turmoil is 0.53.
The proportions of unconditional defectors in groups with bursts (mean = 0.19; s.d. = 0.21) and groups without bursts (mean = 0.49; s.d. = 0.26) are box-plotted in Figure 4 (Welch test $t=4.925;P=4.119$ x $10^{-6}$; df = 54.41; the ROC-curve has AUC = 0.774). The predicted critical threshold (vertical line in the figure) separates the two boxplots, but one might question how robust this result is, given some error in coding. There is a risk of missing a violent act (that can be executed extremely rapidly) or mislabeling a fast movement as violence, thus mixing up fighters and defectors, and a risk of mis-assigning individuals to groups. Even if we suppose there is a 10% chance of each of these errors (which corresponds to Krippendorff’s $\alpha=0.75$; see inter-coder reliability in the Supplementary information), the result is barely weakened (Welch test $t=4.010;P=0.002$, averaged over 1000 simulation runs). Hence, the distinction between bursts and non-bursts is fairly robust.
Figure 4: The proportion of unconditional defectors, and the predicted critical level (vertical line) that separates the boxplots of bursts (1) and non-bursts (0). The data points are our 59 groups.
The empirical distinction between bursts and non-bursts is not perfect, and Figure 4 shows that 14 groups out of 59 (24%) are categorized incorrectly. Most of the errors are due to local circumstances, not a flawed Ising model. In 11 of these 14 groups, the time between the first violent act and last participant joining in was above our two-second interval. Hence, we did not count these collective actions as bursts even though the level of cooperation was higher than expected in non-bursts (Fig 2). One reason for the slow starts could be turmoil lingering at its critical level. In simulations, this lingering causes lags between first instances of violence and the moment when at least half of the group participates (Fig. 5), versus steadily increasing turmoil that yields a proper burst (Fig. 2). However, we would need a better scale of turmoil than the cumulative scale we currently have to assess lingering turmoil. Moreover, in our 11 slow start cases, local circumstances appear to have overridden the Ising dynamics. When re-examining the cases with time lags beyond two seconds, we noticed various causes of delay: (1) group members were in a car and it took time to get out; (2) a focal group member stood at a distance from the fight on a slippery floor next to a pool; (3) a focal group member was beaten and seemed intoxicated by alcohol, hence responded slowly; (4) a focal group member was first pushed over and it took him time to get on his feet again; (5) opponents moved around so fast that it took time to hit one; (6) an opponent fell into thick bushes such that only one focal group member at a time could get to him; (7) group members were sitting in a train and needed time to get on their feet; (8) group members were constrained in a small room where they searched for space to lash out; and, (9) group members were hindered by the mess of two opponents fighting on a collapsed table full of food. One can hardly blame the model for not incorporating these local circumstances that caused delays, and perhaps our two seconds limit should be applied with more lenience in such cases. Finally, two delays were caused by ambiguous behavior: (10) someone tried to de-escalate, was hit, and then attacked, and (11) someone tried to pull back his fighting friend (de-escalation) but almost simultaneously attacked.
Figure 5: Simulation of Equation 1 at approximately the critical level of turmoil, $T_{c}\approx 0.11$, over a range of Monte Carlo steps as proxies for time points. The vertical axis shows the proportion of cooperators ($N_{C}/n$) in a fully connected network ($n=5$) with one unconditional defector (without the unconditional defector, $T_{c}$ is smaller). Note an extended period of multiple short spikes of solitary (and once collective) action before all conditional cooperators join in. Similar lags with spikes occur without unconditional defectors.
In four cases, non-bursts were predicted whereas bursts occurred. Although for parsimony, we applied the same $S_{0}$ (and $\Delta$) for all conditional cooperators, some combatants may actually have a higher value than the average, $S_{0,i}>\overline{S_{0}}$, which could explain these cases.
5.1 Alternative explanation
A plausible alternative explanation for the onset of collective action is synchrony of motion [50], which yields a feeling of oneness among group members [27] and a stronger willingness to take risks for one’s group mates [77].
We measured the synchronization of focal group members pairwise if they engaged in the following behaviors simultaneously (i.e., overlapping intervals on the timeline): approaching members of the opponent group at the same pace (normal, energetic, or running), distancing from the opponent group at normal walking speed (we consider simultaneous distancing at energetic or running pace uncontrolled attempts to escape rather than forms of synchrony), or simultaneously engaging in aggressing. To calculate the level of synchrony, we noted the duration of each instance of synchrony from the beginning of the clip prior to the moment of maximum participation in violence, and we multiplied it by the proportion of focal group ties involved (i.e., normalized with respect to the maximum number of ties in the group).
Although in 21 out of 23 bursts, some degree of synchronization (10.8 on average) preceded collective violence, there were 18 cases in which synchronization (9.6 on average) was not followed by collective violence. In several of the latter cases, synchrony turned out to be a deceptive performance composed of blustering and aggrandizing without commitment to fighting. This does not imply that synchronization is unimportant (it probably is for solidarity; [20]), but it does not predict collective violence.
6 Discussion and Conclusion
The simple Ising model is a century old [13] and has been applied to a wide range of problems [75, 46], to which we add the dilemma of collective action. It explains cooperation parsimoniously, based on agitating stimuli without recourse to strong rationality, initiative takers, norms, feedback through selective incentives, or reliable information passing through the network, resulting in reputations. The Ising model elucidates the temporal unfolding of violence by predicting a critical threshold of unconditional defectors that distinguishes a burst of collective action from a slow start. This pattern is largely supported by the data. The model also explains why violent groups are often small or are small subgroups of larger groups despite greater risk. Small (sub)groups have a lower critical threshold of turmoil, and in a confrontation with opponents, lower levels are reached earlier. We also investigated whether synchronous action precedes violence, but we found that synchronization precedes both collective and solitary violence, and cannot predict either of these outcomes. However, synchronization may still be important to increase solidarity [30, 24].
This study has several limitations. Because we did not select videos without violence, we cannot be certain that violence is caused by turmoil. Moreover, developing a proper scale of turmoil or agitation for field studies is exceedingly difficult, due to variations across domains (e.g., street violence versus grievances about politics), and different effect sizes of different stimuli (e.g., killed family members versus a push that might be quickly forgotten). On top of these challenges, our measurements underestimated turmoil because the videos are left-truncated, and our data depend on camera angle and vision width, inevitably excluding instances of turmoil. Measuring turmoil or agitation is only possible in exceptional situations, for example when protest movements produce it by themselves through accelerated posting of online messages. In these particular cases, the inter-posting time intervals follow a certain pattern (Moore’s law), which makes it possible to predict $T_{c}$ ahead of large street demonstrations [36]. In lab experiments, ethically responsible intergroup conflicts might be examined experimentally, benefiting from controlled circumstances and unambiguous measurement of agitation. Then, causality of turmoil and of the critical level of unconditional defectors might be established or refuted.
Furthermore, there might be error in our coding. If the truth were perfectly known and the coding corrected accordingly, the two boxplots in Figure 4 might shift to a small degree, but, given our robustness check, it is unlikely that this shift would affect our main result.
Another limitation is in the Ising model itself. Despite predicting the threshold of unconditional defectors fairly well, several groups ended up in the non-burst category due to local circumstances (ignored by the model) that caused delays beyond our two second limit. Furthermore, the Ising model does not predict the severity of violence. For future studies, it is important to expand the number and diversity of cases, and to develop better software than is currently available to automatically code videos. Finally, when individuals find themselves more often in similar situations, they will learn, which is easier in smaller groups where they have a larger influence on their payoff [14]. Some will change their strategy, and turn into unconditional defectors [2] who try to exploit other group members and maximize their individual payoff instead of maximizing the group’s payoff. For the model, this would require individuals updating their decision rules at subsequent Monte Carlo steps.
In this first empirical application, we showed that the Ising model can explain the unfolding of violence, and in all likelihood, more discoveries lay ahead. Extensions to norms (as external field $hN_{m}$, with $N_{m}$ norm enforcers) and noise in actors’ information about others’ behavior (i.e., reputations) have been explored in simulations, but in this empirical study, we had no data about norms and reputations, and left them out of our model. In all likelihood, it can be applied to protests [32] and revolts, which break out more often if (rumors say that) a government or its police are weakened [72, 79], analogous to vulnerable individuals in street fights. In protests, where durations are usually counted in days instead of seconds (as our data), it might be easier to distinguish bursts from slow starts than in our small groups, where the distinction sometimes looked gradual. The model also seems applicable to bystanders collectively helping victims under uncertainty [63], and to lynchings [6]. For example, Nussio [57] argued that lynchings in Mexico can be explained through solidarity combined with peer pressure in a network transmitting reputations meshed with local norms. This argument is perfectly consistent with an expanded Ising model, but it lacks the agitation necessary for cooperation to start without centralized organization, as it did empirically. We would argue that the agitation was due to the violation of social norms, of which the lynching victims were accused (e.g., child theft), augmented by the gossip transmitted through the network. In other countries, putative norm violations that agitated crowds and lead to lynchings were rape [12] and blasphemy [3, 65]. The Ising model might even be applied to other species, for example herd bulls defending group members against attacking lions, or quorum sensing bacteria [22].
Taking a broader evolutionary perspective on the role of random noise, multiple studies have found that it can solve coordination problems, such as collective responses to opponents’ threats, by shaking a group loose from its suboptimal (e.g., non-cooperative) state [70], but it can also disturb an optimal state. In the Ising model, both can happen, depending on the amount of turmoil: when in a group, turmoil approaches $T_{c}$ from below, random noise in decision making facilitates cooperation, but at high levels, too much noise entails confusion and error.
—Supplementary Material—
7 Data and compilation of data set
In contrast to prior studies of violence and deescalatory action based on CCTV footage [42, 43, 63] most of our videos were captured on mobile phones, which tend to provide better picture clarity, detail and sound. Whereas CCTV footage is most often shot from an elevated and fixed camera angle, people who record violent incidents on their mobile phones tend to follow the action, thereby allowing us to observe bodily positioning and movements in more detail [44].
One concern about phone-recorded clips is that recorders start filming when the antagonism is already ongoing; our data are thus left-truncated to an unknown degree. We discarded footage that immediately started with physical violence. Another concern is that uploaded videos might be biased toward more spectacular cases. However, our sample shows variety in the forms and severity of violence, ranging from groups who engage in frenzied collective violence to incidents that involve just a few slaps. A third concern is whether the recording influenced the behavior of the assailants. Several videos showed more than one person recording the incident. A host of research indicates that people do not change their behavior substantially in the presence of a camera and mainly focus on what they are doing (reviewed by [38]). This is probably even more the case in antagonistic situations, and because young people are used to being filmed by phones. Additionally, when people do pay attention to the camera, which happened in just one of our clips, it is visible in the recording and can be incorporated in the analysis. (In our case, this concerned behavior coded as turmoil.)
We used the following criteria to compile our dataset. First, we only included incidents in which at least one of the antagonistic parties comprised at least two members. Second, we only selected clips in which violence could be observed from the (or a) beginning until the end. Third, we focused on fights that seemed to have occurred spontaneously, excluding prearranged fights on the basis of any of the following criteria: the fighters and bystanders all had approximately the same young age (e.g., school yard fights); the antagonists were wearing protective clothing; or a referee was present. Regarding gender, the overwhelming majority of the incidents were between males, with only 8 cases involving both females and males. We excluded female-only fights because they often seemed being arranged (back yard fights with male referees).
Along these criteria, we obtained 42 videos from websites such as YouTube, LiveLeak, and WorldStarHipHop, using search terms with the English keywords “brawl,” “street fight,” and “assault.” This sample appears to be random with respect to temporal unfolding and (sub)group size. 36 clips are from English-speaking countries (mainly the US and the UK, with one from Canada and one from India); 5 of the remaining clips are from the Netherlands, and 1 is from Colombia. We did not observe differences in relevant behavior related to the location of the recording.
The shortest clip lasted 30 sec. and the longest nearly 5 minutes (mean 101 sec.; s.d. 59 sec.). Out of a potential 2 x 42 groups, where the opponents in one analysis become the focal group in the next, 25 groups fought with a single individual rather than a group, which leaves $\mathcal{N}$=59 groups to examine. Most groups were small, $2\leq n<10$ (mean 3.6), but one had 14 members (of which 6 became violent). The smaller ones were simulated as fully connected networks wherein everyone could see one another unless there were obstacles or deescalators obstructing visual contact. Obstruction was simulated by removing $m$ ties.
8 Coding
We used Noldus Observer XT 14 software to code the behaviors of individuals, resulting in start times and durations of behaviors per individual throughout the video clip. The software distinguishes state events of relatively longer duration (e.g., holding a person) from point events that are brief single acts (e.g., a punch). When an act of violence took a more extended duration (e.g., holding or dragging), we coded the entire duration of the act, thus creating a state event. When a series of violent point events appeared less than two seconds after one another, we connected them, thus creating a state event.
To enhance the precision of our coding, we coded observable actions (such as kicking, punching, pointing at opponent) and later lumped them together to generate the main codes: violence, turmoil, and deescalation. We coded the behavior of each actor in multiple rounds, coding only one type of action per round. The video material was played at half speed and repeated many times, which is important because violence proceeds too fast to see what is going on in real time. The first 20 clips were coded by one of the authors and two assistants. We discussed the coding when we disagreed and continued the procedure by multiple coders until agreement was reached. At this point, the remainder of the clips were coded by one (meanwhile) experienced assistant researcher and were later checked by another. The remaining disagreements or mistakes were discussed, and the coding was adjusted accordingly.
To identify whether individuals belonged to a focal, opponent, or third-party group, we observed whether they remained close (at touching distance) to each other as they moved through space, whether they called each other by their names, whether they touched each other (for instance, by tapping shoulders), or whether they shared a vehicle or other object (for instance, a bag or skateboard). One of the authors and two assistants discussed each case in which we observed one or more of these behaviors to determine whether the actors were part of the same group. To assess the reliability of our coding, we asked three students who were unaware of our theoretical expectations and preliminary results to code group memberships, maximum number of participants involved in violence, and the occurrence of bursts in 37 videos, containing 54 groups and 20 solitary individuals (according to our own coding). Notice that just like biologists observing primates, much experience is required to become an accurate observer, which these students did not have. Krippendorff’s alphas for group membership, fighters, and bursts were .87, .68, and .75 respectively, or .77 on average; in the main text we reported that an error percentage of 10% results in .75. We also calculated whether the distinction between bursts and fizzles in students’ coding followed the predicted critical proportion of unconditional defectors. We arrived at the number of unconditional defectors (which the students did not code) by deducing the maximum number of participants in violence from the group size. Out of 49 groups (5 cases missing or coded as individual rather than group, as we did) students identified 29 bursts and 20 fizzles (we required that at least 2 out of 3 students agreed on the observation of a burst to denote it as such). The proportion of defectors was .18 (sd =.03=.03) in bursts and .33 (sd =.04=.04) in fizzles (Welch test 8.3, df 35.6, $P=.007$), with bursts clearly below and fizzles right at our critical threshold.
9 Composition of indicators
Violence. We coded violence when members of the focal group used force against another’s person’s body (by punching, slapping, kicking, hitting, stomping) and/or when actors moved or held another person’s body forcefully (by pushing, shoving, dragging, wrestling, holding, etc.). To enhance the reliability of our coding, we also coded whether actors used their hands, legs, or other body parts (e.g., elbow, head) and whether weapons were used, including the type of weapon, excluding guns (i.e., knives, sticks, tasers and improvised weapons, such as a skateboard or bottle, etc.). We define collective violence as time intervals wherein two or more members of the focal group used violence simultaneously, either as overlapping state events or, in fewer cases, as simultaneously occurring point events. The maximum participation in collective violence was determined by taking the largest number of focal group members engaging in violent action, divided by the total number of focal group members.
In the videos, it was not possible to distinguish leaders from initiative takers, but we noticed individuals who started violence on their own. In simulations, a leader/initiative taker $i$ can be incorporated by a larger, individual $S_{0,i}$ value; the result is collective action at lower $T$.
Unconditional defectors do not cooperate when $T\geq T_{c}$ or most network neighbors cooperate. First, we labeled as unconditional defectors, the focal group members who took deescalatory action towards others. Their behavior was coded as follows [42]: open-handed gestures in the direction of other individuals; waving arms to stop or dampen in the direction of others; touching or patting; guiding a person away; pulling people apart; and putting one’s body in between opponents. Whether deescalators’ interventions were effective or not, they were at least temporarily unavailable to participate in violence themselves, and their behavior signaled noncooperation to their fellow group members. Second, we noted group members who were effectively stopped from using violence for at least five seconds by other focal group members, opponents, or third-party members (e.g., bystanders). Third, group members remaining passive when others fought. Fourth, group members unable to participate in violence due to spatial constraints for at least five seconds, for instance, due to cars blocking their way or because they were not close enough to the action. Fifth, group members unable to participate for at least five seconds because they were wounded or had fallen to the ground. We took all five reasons for and causes of defection into account for the proportion of unconditional defectors before and during the first instance of violence by the focal group.
Turmoil. We subsumed the following behaviors of members of the opponent group under the heading of turmoil for the focal group. First, we coded approaching when opponents moved closer to focal group members. We also indicated whether actors moved energetically (jumping/skipping) and whether they ran. Second, aggressive behaviors when actors in the opponents’ group moved body parts in a way that signaled provocation, hostility and/or a readiness to attack. We used the following modifiers (sub-codes) to specify aggressing behavior: fighting gestures; pulling off clothing (jackets or vests); pulling up pants; pointing at opponents; provocative gesturing with fingers or hands (as an invitation to engage); bending forward (head and/or chest bending toward opponent); encroaching (invading opponents’ personal space through using or damaging objects belonging to them); and teasing (invading opponents’ body space in a teasing/ridiculing way that does not inflict bodily harm, such as lightly hitting, stroking, or patting the opponent’s body). Third, involuntary behavior that increases the vulnerability of opponents, because it tends to agitate (signaling opportunity) and provoke violence. This concerned stumbling and falling to the ground. Finally, violence committed by members of the opponent group. Because we do not know the psychological effect sizes of the various kinds of turmoil, we could not construct a ratio level scale. Instead, we calculated the total level of turmoil from the beginning of the clip until a focal group’s maximum participation in violence by noting the duration of each instance of turmoil, and multiplying it by the number of opponents involved, i.e., the surface area under the turmoil curve in Fig. S9, S12, and S13.
Bursts. We identified bursts when, at the first moment of collective violence, two conditions are fulfilled: (1) at least half of the focal group members joined the fight, or both actors did in a dyad, and (2) cooperators started fighting in less than 2 seconds (with a 5% margin) after the first. Condition (1) follows from simulations of small groups; see Table S3. Note that in simulations of much larger networks ($n\geq 1000$), bursts can occur wherein less than half of a group participates. In some videos, we observed multiple bursts; in these cases, we only considered the first burst to facilitate comparison of cases.
10 Examples from videos
Graphs of behavior on the timeline enable to analyze and compare the coded videos. Each graph shows the intervals and their durations of (collective) violence, turmoil, synchrony and deescalation (see Fig. S9, S12, S13).
To illustrate, we discuss how coded video data was turned into a graph for clip 26. Fig. S6-8 feature screenshots of the clip, and Fig. S9 displays the graph. This clip features a confrontation of a dyad (focal group) with a triad (opponents) in a restaurant. Violence started at approximately 14 seconds into the video and quickly became collective until approximately 28 seconds. Four instances of turmoil preceded the violence, starting with one of the members of the opponent group gesturing aggressively and walking around in a jumpy, energetic way. The members of the focal dyad walked synchronously toward and away from the opponent in three shorter and one relatively longer stretch of time. Fig. S6 captures a moment of synchrony; the members of the dyad were making fighting gestures and stood aligned, each stretching out one leg with their feet pointing toward the opponent, and the other leg standing backward. Fig. S7 shows the moment when the dyad’s level of agitation reached a critical level; when one of the focal actors dodged backward, avoiding being hit, his companion was about to deliver a blow to the head of the opponent, which made the latter fall down. Fig. S8 displays the second member of the opponents’ triad (soon joined by the third) rushing into the restaurant toward the focal dyad, near his fallen companion. At this point, both groups were engaged in collective violence. Fig. S7 also shows bystanders (who can be seen sitting at a table in Fig. S6) and employees standing near the counter. At 28 seconds, an employee took deescalatory action for the first time. At that point, the collective violence had stopped already and erupted again for a shorter moment approximately 2 seconds later. The clip ends when two triad members carry their fallen group member outside. By then, the focal dyad had already left the scene.
Figure 6: Still taken from clip 26 at 0:06 seconds.
Figure 7: Still taken from clip 26 at 0:13 seconds.
Figure 8: Still taken from clip 26 at 0:15 seconds.
Figure 9: Graph of clip 26 with a burst of collective violence by a dyad.
Fig. S10-S12 provide another illustration. Fig. S12 is the graph of clip 86; it shows three outbreaks of collective violence in a group of five members who faced a single opponent. In the first instance of collective violence (not a burst by our 2 seconds criterion), starting at approximately 60 seconds into the clip and lasting for approximately 30 seconds, four members participated. Three of them immediately followed each other, and the fourth joined after approximately 2 seconds. Prior to the outbreak of violence, the opponent agitated focal group members by walking back and forth and by posturing aggressively. Fig. S10 captures the trigger moment of agitation; i.e., while one member of the focal group was engaged in aggressive posturing and a group member tried to deescalate, the opponent was about to strike another focal group member. Fig. S11 shows the situation 4 seconds later, when three members of the focal group were attacking.
Figure 10: Still taken from clip 86 at 0:59 seconds.
Figure 11: Still taken from clip 86 at 1:03 seconds.
Figure 12: Graph of clip 86. Outbreaks of collective violence in a group of five. Because the time between the first and last participant’s commencement of violence exceeds 2 seconds, these outbreaks are not classified as bursts.
11 Analysis of coded and plotted videos
We used graphs of coded videos in our analysis, illustrated in Fig. S13, where the opponent group in clip 67 is plotted. The horizontal axis is the timeline in thousands of seconds, the vertical axis is the proportion of group members involved (for turmoil and violence), and the proportion of intragroup ties involved out of the maximum possible number of ties (for synchrony of action). The first moment of collective violence, denoted by the downward arrow, involves two focal group members. There is, however, a short interruption of collective violent action. Therefore, we did not code this occurrence of collective violence as a burst. The time gap between the first participant and maximum participation in collective violence, denoted by the horizontal arrows, takes over 6 seconds.
In the figure, seven instances of turmoil occur (numbered 1-7) prior to the moment of maximum participation in violence: (1) a short interval that, after a brief interruption, continues for 4 seconds and involves 67% of the opponent group (2 members); (2) involves all three members of the opponent group and lasts nearly 1 second; (3) lasts less than 1 second and involves 33% of the opponent group; (4) turmoil continues, with an intermittent short dip, for over 1 second with 67% involvement; (5) a peak at 100% involvement that lasts less than 1 second; (6) a short drop to 67% involvement that takes approximately 1 second; (7) the last turmoil before maximal violence. We calculated the level of turmoil by multiplying the time length of each of these intervals by the number of actors involved.
Intervals of synchrony in Fig. S13 are indicated by A and B. A lasted 2.4 seconds and involved all ties in the group; B involved 67% of the ties for approximately 2 seconds prior to maximum participation in violence. Finally, the graph shows intervals of deescalatory action.
Figure 13: Graph of clip 67; opponent group. The timeline displays the start of collective violent action (downward arrow), instances of turmoil (1-7), synchrony (A,B), the time lag between the first and last group member joining the collective violence (horizontal arrows), and deescalatory action.
12 Overview of cases and descriptive statistics
Our dataset contains information about the behavior of 406 individuals in 42 violent incidents. The 59 groups and their key characteristics are listed in Table SLABEL:table:long. After focal and opponent groups have been analyzed as listed, they flip roles with respect to cooperation and turmoil production in a second batch of the analysis, except when an opponent is solitary instead of a group.
group $n$ c.v. $N_{C}$ lag burst $T$ Sync $N_{D}$
1 1_beachfight_focal 3 1 2 0.00 1 0.97 19.39 1
2 3_asianeateries_focal 5 0 1 0 72.41 18.10 4
3 3_asianeateries_oppon. 2 0 1 0 93.37 16.20 1
4 6_rikshadrivers_focal 6 1 4 1.93 1 7.24 5.33 2
5 7_baldies_focal 4 1 3 0.92 1 6.42 0.94 1
6 8_queensday_focal 14 1 6 1.95 1 45.81 303.42 8
7 9_gasstation_focal 2 0 1 0 14.49 16.30 1
8 10_redpickup_focal 3 1 3 0.93 1 9.34 78.75 0
9 26_insideeatery_focal 3 1 3 1.21 1 14.41 0.20 0
10 26_inseeatery_oppon. 2 1 2 1.77 1 7.49 5.32 0
11 27_dallasairport_focal 6 1 5 0.00 1 24.99 3.53 1
12 32_shovedbycar_focal 3 1 2 3.66 0 3.66 0.00 1
13 41_kababshop_focal 2 0 1 0 42.12 0.00 1
14 41_kababshop_oppon. 2 1 2 0.84 1 5.90 0.42 0
15 45_onbusystreet_focal 2 1 2 1.48 1 12.79 9.35 0
16 50_fitness_focal 2 1 2 3.08 0 4.62 1.54 0
17 52_blockingskaters_focal 2 0 1 0 20.44 10.72 1
18 52_blockingskaters_oppon. 4 0 1 0 45.56 17.36 3
19 53_famousonyoutube_focal 2 1 2 20.23 0 14.45 0.00 1
20 53_famousonyoutube_opp. 2 0 1 0 51.11 0.00 1
21 54_hammeronhead_focal 4 1 3 1.57 1 9.17 4.19 1
22 54_hammeronhead_oppon. 2 1 2 1.57 1 31.43 5.24 0
23 60_gangattack_focal 6 0 1 0 31.73 80.43 5
24 61_burgerking_focal 2 0 1 0 37.88 0.00 1
25 66_guyvsgang_focal 6 1 6 1.81 1 0.91 19.22 0
26 67_poolfight_focal 3 1 2 0.84 1 16.46 4.36 1
27 67_poolfight_oppon. 3 1 3 6.29 0 19.87 2.49 1
28 84_brokenwindow_focal 3 0 1 0 35.36 2.89 2
29 86_tables_focal 5 1 4 2.25 0 29.23 4.50 1
30 88_babypowder_focal 2 1 2 2.10 1 5.43 4.20 0
31 92_didntdoanything_focal 2 0 1 0 46.51 8.35 1
32 93_groupkicking_focal 4 1 3 5.92 0 8.68 10.02 1
33 95_apologise_focal 9 1 2 1.51 0 26.42 76.30 7
34 95_apologise_oppon. 2 0 1 0 133.06 2.64 1
35 96_gangversusbat_focal 6 0 1 0 25.00 67.06 5
36 97_alleyfight_focal 4 0 1 0 61.14 26.32 4
37 97_alleyfight_oppon. 4 1 2 0.76 1 34.39 3.83 2
38 98_stopmick_focal 3 1 2 0.76 1 20.61 17.54 1
39 98_stopmick_oppon. 2 0 1 0 68.10 10.70 1
40 101_fightandrob_focal 6 1 4 2.97 0 22.19 21.45 2
41 101_fightandrob_oppon. 2 0 1 0 79.09 3.10 1
42 104_waiting_focal 4 1 4 0.73 1 13.83 0.36 0
43 104_waiting_oppon. 2 0 1 0 128.82 0.36 1
44 105_frenziedassault_focal 3 1 2 1.21 1 0.00 9.61 1
45 108_sidewalkbrawl_focal 2 0 1 0 268.40 44.73 1
46 108_sidewalkbrawl_oppon. 4 1 2 0.00 1 33.55 36.72 2
47 110_nietklaar_focal 5 1 3 14.45 0 36.13 12.04 2
48 121_eightballjacket_focal 4 1 3 2.66 0 8.64 1.08 1
49 124_leavehimalone_focal 3 1 3 4.86 0 13.36 0.00 0
50 124_leavehimalone_oppon. 2 1 2 0.00 1 9.09 0.00 0
51 128_throwingfood_focal 2 1 2 1.14 1 0.57 0.00 0
52 128_throwingfood_oppon. 2 1 2 2.28 0 6.83 0.00 0
53 129_awayfrommyfood_focal 5 1 2 3.06 0 14.22 4.09 3
54 129_awayfrommyfood_oppon. 6 0 1 0 65.37 30.00 5
55 133_homeboy_focal 2 1 2 2.68 0 0.54 1.07 0
56 134_down_focalgroup 4 1 2 0.89 1 23.12 45.27 2
57 134_down_oppon. 5 1 2 0.89 0 40.02 9.34 3
58 138_cometomyhouse_focal 3 0 0 0 34.02 36.51 3
59 145_knifefightcolombia_focal 2 0 1 0 17.04 20.16 1
Table SLABEL:table:long: Overview of the groups. Occurrence of collective violence (c.v.); maximum participation in violence ($N_{C}$); time lag between the first participant and maximum participation in collective violence; burst; levels of opponents-produced turmoil ($T$) and of synchrony at the first outbreak of c.v.; and, number of unconditional defectors ($N_{D}$). In case 6, two focal group members disappeared from view before the first moment of violence had started; out of the remaining 12, 6 participated in collective violence. In cases 19, 27 and 36, $NC+ND\neq n$ due to participants’ role switching. In cases 19 and 27, one of the group members defected first (deescalated; was not present at the start) but then switched to fighting. In case 36, the single fighter fell to the ground.
Table S2 provides mean values, percentages, standard deviations, and minimum and maximum values for the indicators used in the analysis.
mean or $n$ (%) s.d. min max
1 Group size 3.58 2.09 2 14
2 Dyads 24 (40.7) 0 1
3 Triads 11 (18.6) 0 1
4 Groups $\geq 4$ members 24 (40.7) 0 1
5 Collective violence 38 (64.4) 0 1
6 Burst 23 (39) 0 1
7 Maximum participation in violence .64 .28 0 1
8 Average time lag in bursts in sec. ($n=23$) 1.06 .65 0 2.10
9 Average time lag per participant in non-bursts ($n=15$) 5.12 5.29 0.89 20.23
10 Turmoil prior to max participation in violence ($n=38$) 37 (97.4) 0 1
11 Level of turmoil prior to max part. in violence ($n=38$) 15.34 12.21 0 45.81
12 Synchrony prior to max part. in violence ($n=38$) 32 (84.2) 0 1
13 Level of synchrony prior to max part. in violence ($n=38$) 18.98 50.95 0 303.42
14 Proportion unconditional defectors .37 .28 0 1.0
Table 2: Descriptive statistics. Note: Levels of synchrony prior to maximum violence includes an outlier of 303.42. Without outlier, mean = 11.29; s.d. = 18.96; max = 78.75. We excluded this outlier in the calculations reported.
13 Solving the Ising model
The Ising model can be solved computationally by the Metropolis algorithm (main text). Here, it is solved analytically by a mean field analysis, first for $p=0$ and then $p>0$. Subsequently, deviations of the computational results from the mean field result are pointed out.
13.1 Mean field analysis
For the mean field analysis, we assess the level of cooperation, $N_{C}/n$, in terms of the order parameter, $M=1/n\sum_{i=1}^{n}S_{i}$. Consequently, $N_{C}/n=(M+D)/(C+D)$. The mean field assumption can be stated as $S_{i}=\bar{S_{i}}=M$. For the moment, the proportion of unconditional defectors, $p=0$; see below when $p>0$.
We start out with the Hamiltonian, $H=-\sum_{i,j}w_{ij}S_{i}S_{j}$. We use the mapping $\{C,-D\}\rightarrow\{S_{0}+\Delta,S_{0}-\Delta\}$ with bias $S_{0}=(C-D)/2$ and offset $\Delta=(C+D)/2$ to rewrite the Hamiltonian as
[…]
To calculate the Boltzmann probabilities of a single spin (or an individual’s probabilities to cooperate or defect), we define the pertaining Hamiltonian
[…]
$H_{i}$ $=-\sum_{j}w_{ij}(S_{0}+\hat{S_{i}})M$
$=-(S_{0}+\hat{S_{i}})M$
$H_{i}^{\pm}$ $=-S_{0}M\mp\Delta M.$ (6)
In the subsequent derivation, we use no Boltzmann constant (and no $\beta$ either), just $T$. The average value of a spin, $\bar{S_{i}}$, according to the Boltzmann distribution, with $P(S_{i}^{-})$ standing for the probability that $S_{i}$ is negative and $P(S_{i}^{+})$ that it is positive, is
$\bar{S_{i}}$ $=S^{-}P(S_{i}^{-})+S^{+}P(S_{i}^{+})$ (7)
$=\dfrac{S^{-}e^{-H_{i}^{-}/T}+S^{+}e^{-H_{i}^{+}/T}}{e^{-H_{i}^{-}/T}+e^{-H_{i}^{+}/T}}$
$=\dfrac{S^{-}e^{-(-S_{0}M+\Delta M)/T}+S^{+}e^{-(-S_{0}M-\Delta M)/T}}{e^{(-S_{0}M+\Delta M)/T}+e^{-(-S_{0}M-\Delta M)/T}}$
$=\dfrac{S^{-}e^{-\Delta M/T}+S^{+}e^{\Delta M/T}}{e^{-\Delta M/T}+e^{\Delta M/T}}$
$=\dfrac{(S_{0}-\Delta)e^{-\Delta M/T}+(S_{0}+\Delta)e^{\Delta M/T}}{e^{-\Delta M/T}+e^{\Delta M/T}}$
$=S_{0}\dfrac{e^{-\Delta M/T}+e^{\Delta M/T}}{e^{-\Delta M/T}+e^{\Delta M/T}}+\Delta\dfrac{-e^{-\Delta M/T}+e^{\Delta M/T}}{e^{-\Delta M/T}+e^{\Delta M/T}}$
$=S_{0}+\Delta\tanh{(\Delta M/T)}.$ (8)
This result without unconditional defectors was inferred in an earlier study (author(s)). Here, we also deal with unconditional defectors in proportion $p>0$. Accordingly, we define $M_{cc}$ as the average spin value of the conditional cooperators and $M_{ud}$ as the average spin value of the unconditional defectors. Note that $M_{ud}=S^{-}$. We assume that the unconditional defectors are homogeneously distributed across the network. Accordingly, the mean field equation becomes
$S_{i}$ $=pS^{-}+(1-p)M_{cc}.$ (9)
The Hamiltonian for a single conditional cooperator becomes
$H_{i}^{\pm}$ $=-(S_{0}\pm\Delta)(pS^{-}+(1-p)M_{cc})$ (10)
$=-S_{0}pS^{-}\mp\Delta pS^{-}-S_{0}(1-p)M_{cc}\mp\Delta(1-p)M_{cc}.$ (11)
In the derivation of Eq. 8, all terms that did not contain $\mp\Delta$ canceled each other out. For clarity, we remove these terms from Eq. 11, which results in
$H_{i}^{\pm}$ $=\mp\Delta(pS^{-}+(1-p)M_{cc}).$ (12)
The mean field analysis for conditional cooperator $i$ is
$\bar{S_{i}}$ $=S^{-}P(S_{i}^{-})+S^{+}P(S_{i}^{+})$
$=\dfrac{S^{-}e^{-\Delta(pS^{-}+(1-p)M_{cc})/T}+S^{+}e^{\Delta(pS^{-}+(1-p)M_{cc})/T}}{e^{-\Delta(pS^{-}+(1-p)M_{cc})/T}+e^{\Delta(pS^{-}+(1-p)M_{cc})/T}}$
$=S_{0}+\Delta\dfrac{-e^{-\Delta(pS^{-}+(1-p)M_{cc})/T}+e^{\Delta(pS^{-}+(1-p)M_{cc})/T}}{e^{-\Delta(pS^{-}+(1-p)M_{cc})/T}+e^{\Delta(pS^{-}+(1-p)M_{cc})/T}}$
$=S_{0}+\Delta\tanh{(\Delta(pS^{-}+(1-p)M_{cc})/T)}=M_{cc}.$ (13)
Using Eq. 9, we can express the self-consistency equation of $M_{cc}$ in $M$,
$M$ $=pS^{-}+(1-p)(S_{0}+\Delta\tanh{(\Delta(pS^{-}+(1-p)M_{cc})/T)})$
$=pS^{-}+(1-p)(S_{0}+\Delta\tanh{(\Delta M/T)})$
$=S_{0}-p\Delta+(1-p)\Delta\tanh{(\Delta M/T)}.$ (14)
13.2 Critical proportion of unconditional defectors
Depending on $T$, the self-consistency equation has one unstable and two stable ferromagnetic solutions, or one stable paramagnetic solution. At a critical $T$, the system transitions between these two states. When the system is paramagnetic,
$\frac{\partial}{\partial M}(S_{0}-p\Delta+(1-p)\Delta\tanh{(\Delta M/T)})<1$
at the solution of $M$. When the system is ferromagnetic,
$\frac{\partial}{\partial M}(S_{0}-p\Delta+(1-p)\Delta\tanh{(\Delta M/T)})>1$
at the unstable solution of $M$. We can identify a critical $T$ when
$\frac{\partial}{\partial M}(S_{0}-p\Delta+(1-p)\Delta\tanh{(\Delta M/T)})=1,$ (15)
hence,
$\frac{1}{\cosh^{2}{(\Delta M/T)}}\Delta^{2}(1-p)/T$ $=1$
$\cosh{(\Delta M/T)}$ $=\sqrt{\Delta^{2}(1-p)/T}$
$\Delta M/T$ $=\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}$
$\frac{M}{\Delta}$ $=\frac{\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}}{\Delta^{2}/T}.$ (16)
We can substitute this expression (16) in the self-consistency equation (14)
$\frac{\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}}{\Delta/T}$ $=S_{0}-p\Delta+(1-p)\Delta\tanh{({\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}})}$
$\frac{\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}}{\Delta^{2}/T}$ $=\frac{S_{0}}{\Delta}-p\pm(1-p)\tanh{({\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}})}$
$=\frac{S_{0}}{\Delta}-p\pm(1-p)\sqrt{1-\frac{1}{\Delta^{2}(1-p)/T}}.$ (17)
Discarding the equation that has no real numerical solutions leaves
$\frac{\operatorname{arcosh}(\sqrt{\Delta^{2}(1-p)/T})}{\Delta^{2}/T}+(1-p)\sqrt{1-\frac{1}{\Delta^{2}(1-p)/T}}=\frac{S_{0}}{\Delta}-p.$ (18)
Solutions of this equation become complex if $p>\frac{S_{0}}{\Delta}$, hence $p\leq\frac{S_{0}}{\Delta}$. Graphical illustrations are in Fig. S14. The choice of $C=1$ and $D=1/2$ implies that there is no (burst of) cooperation if $p>1/3$.
Figure 14: Given $S_{0}=0.2$, the critical level of agitation, $T_{c}$, is plotted as a function of the proportion of unconditional defectors, $p$, for various levels of $\Delta$. The critical level, $p_{c}$, is reached at the right-hand end of the lines.
13.3 Solving the Ising model computationally
The Ising model was solved computationally, in Fortran for speed and in R for comfort. Surprisingly, $p_{c}$ of simulated, fully connected networks is very close to the mean field calculation of $p_{c}$, even in very small networks. Given $C=1$ and $D=1/2$, there is no burst if $p_{c}>0.34$. The critical threshold is nearly density independent: if in (sufficiently large) simulated networks, density is decreased by two orders of magnitude, $p_{c}$ increases from $\approx 0.34$ to $\approx 0.35$. The degree distribution has no effect on $p_{c}$.
In simulations, $p_{c}$ is less precise in smaller networks (Table S3) and the prediction for the triad is empirically false, whereas the mean field is correct: one unconditional defector does not prevent cooperation of the other two. Also one empirical group of six with two unconditional defectors had a burst, which did not happen in the simulations (Fig 15).
$n$ $N_{D}$ $\approx N_{C}$ burst
2 1 0 0
3 1 1 0
4 1 2 1
5 1 3-4 1
6 1 4-5 1
5 2 1 0
6 2 2 0
7 2 4 1
8 2 4-5 1
9 2 6 1
9 3 3 0
Table 3: Simulations just below (burst = 1) and above (burst = 0) the critical threshold of unconditional defectors ($N_{D}$) in small groups (size $n$). The numbers of cooperators ($N_{C}$) fluctuate across simulation runs and are approximate. Note: For groups with $n=7$ and $N_{D}=2$, simulations in Fortran yield $N_{C}\approx 4$ whereas in $R$, $N_{C}\approx 3$.
Figure 15: Color online. For simulated networks with sizes $3\leq n\leq 100$, the minimum number of unconditional defectors who prevent a burst is plotted. The red dot (left, closest to the bottom) marks the triad (1 defector prevents a burst of 2), which conflicts with the empirical data (1 defector and a burst of 2).
Authors’ contributions
JB made the asymmetric Ising model, and wrote the software and the paper. DW collected, interpreted, and analyzed the data. BM did the mean field analysis. JB and DW wrote the Supplementary Material. No AI has been used in this research project.
Data availability
All coded video data, the R script used to produce plots from the data (Supplementary information), and a Fortran script for simulations of the Ising model are available at https://osf.io/f25nq/